OpenStudy (anonymous):

f(x) is polynomial (x>0) f(x)+f(1/x)=f(x)f(1/x) and f(2)=-7 find f(3)?

OpenStudy (perl):

sorry the english is confusing me. 'now the other one f(x)' , what other one ?

Parth (parthkohli):

I think he means "another question". A nice one though.

OpenStudy (anonymous):

yes it is another question and i edited the question

OpenStudy (anonymous):

@iGreen

Parth (parthkohli):

@Zarkon @Callisto

OpenStudy (perl):

So the conditions are f(x)+f(1/x)=f(x)*f(1/x) and f(2)=-7

OpenStudy (perl):

what kind of function has this property f(x) + f(y) = f(x) * f(y)

OpenStudy (anonymous):

here is the tip f(x) =1+x^n or 1-x^n

OpenStudy (anonymous):